# EvGGS: A Collaborative Learning Framework for Event-based Generalizable Gaussian Splatting

### v1

@@ Abstract
Event cameras offer promising advantages such as high dynamic range and low latency, making them well-suited for challenging lighting conditions and fast-moving scenarios. However, reconstructing 3D scenes from raw event streams is difficult because event data is sparse and does not carry absolute color information. To release its potential in 3D reconstruction, we propose the first event-based generalizable 3D reconstruction framework, called EvGGS, which reconstructs scenes as 3D Gaussians from only event input in a feedforward manner and can generalize to unseen cases without any retraining. This framework includes a depth estimation module, an intensity reconstruction module, and a Gaussian regression module. These submodules connect in a cascading manner, and we collaboratively train them with a designed joint loss to make them mutually promote. To facilitate related studies, we build a novel event-based 3D dataset with various material objects and calibrated labels of grayscale images, depth maps, camera poses, and silhouettes. Experiments show models that have jointly trained significantly outperform those trained individually. Our approach performs better than all baselines in reconstruction quality, and depth/intensity predictions with satisfactory rendering speed.
@@ 1 Introduction
The contributions of this paper are summarized as follows: (1) We first propose the pure event-based, generalizable 3DGS framework (EvGGS), which faithfully reconstructs 3D scenes as 3D Gaussians from raw event streams and generalizes to various unseen scenarios. The proposed method outperforms existing event-based methods by a large margin.
@@ 5.1 Event-based 3D dataset
Dataset Existing event-based 3D datasets such as (Rudnev et al., 2023; Zhou et al., 2018) only contain a limited number of objects and lack high-quality intensity, depth, and mask groundtruths because they mainly concentrate on single scene reconstruction or sparse vision tasks. To fill the current gaps in the community, we establish a full event-based 3D dataset including completed labels, referred to as Ev3DS. The dataset includes a wide variety range of materials. There are 64 objects for training and 15 objects The dataset is constructed and rendered via Blender, it encompasses a multitude of photo-realistic objects, characterized by their complex and varied geometric structures and texture information. We have harnessed VisionBlender(Cartucho et al., 2020) to gather dense depth, mask, and pose information. In each scene, events are generated by a virtual event camera that orbits around the origin of the object in space. We employ V2E(Hu et al., 2021) to generate synthetic event streams maintaining default noise configurations. Additionally, to verify the robutness of the proposed method, we establish and release a novel realistic dataset utilizing the event camera DVXplore for further faithful evaluations, referred to as Ev3D-R. The real-world dataset is essential because the real event camera will raise noises and be more sensitive to illumination changes than synthetic event data. We evaluate the scalability and generalization of the proposed methods on the realistic event data. Detailed introduction of Ev3D-R can be found in the Appendix.
@@ 5.2 Performance of neural reconstruction
Table 1 reports all metrics to quantitatively compare the results, which illustrates that our method achieves the best performance. Gen. means whether the corresponding methods can generalize to unobserved scenes. Real-time means the corresponding methods can render several images within 1 second. More analysis of this point is in the Appendix. Figure 2 presents the qualitative results.
@@ 5.4 Quality of depth estimation
5.5 Quality of realistic event data
@@ 5.6 Ablation studies
This subsection demonstrates the impact of different training strategies on model performance in the tasks of novel view synthesis and depth estimation. w/o Joint denotes that we only train the Gaussian regressor with the individually trained and frozen depth estimator and intensity reconstructor. w/o Cascade means that the input does not contain the feature map of the previous network, but only the event voxel and prediction results from the last modules. In Table. 4, the performance of w/o Joint and w/o Cascade is significantly degraded because submodules hardly benefit from the others in the two settings. Besides, w/o ${L}_{I}$ and w/o ${L}_{D}$ represent the corresponding loss variants by removing $L_{I_{\theta}}$ and $L_{D_{\phi}}$ respectively. Table. 4 demonstrates that the absence of depth supervision during joint training leads to a decline in depth estimation. Additionally, w/o $L_{I}$ results in a lack of constraints for intensity reconstruction, which degrades the performance of the subsequent cascaded Gaussian regressor and adversely affects the other two submodules with varying degrees.
@@ 6 Conclusion
7 Impact statements
This work aims to advance the pure event-based 3D reconstruction to release the potential of the brain-inspired camera including high dynamic range, low temporal latency, and low energy consumption, etc. This work is the first to use only event data to reconstruct 3D Gaussians in a generalizable way.
@@ Appendix F Additional Experiments
F.1 Realistic Dataset
In this subsection, We give a detailed introduction to our data collection pipeline. This realistic dataset is captured by the DVXplore event camera. First, we render RGB images and generate corresponding depth groundtruth via Blender, this step is similar to the synthetic dataset. Then we display the videos for these objects on a high fresh rate screen in our work studio with constant low lighting conditions. Meanwhile, we use the DVXplore to continuously capture the screen to obtain the corresponding realistic event stream. Before collecting data, we have already well-calibrated these devices to ensure high-level data association by using chessboard calibration and image, depth, and event frame alignment. By using the data collection pipeline, we can obtain a realistic event dataset with corresponding images, depths, and 3D model labels as well. Moreover, The distribution of events captured by the real DVXplore will be more complex, realistic, and disordered, which places higher demands on models.
@@ F.2 Qualitative Experiments on Large Scale Scenes
We compare our method with GrayNeRF and E-ENeRF. GrayNeRF means we trained the original NeRF by the grayscale images converted from the RGB images. We optimize the GrayNeRF from scratch because it is a per-scene optimization method. For E-ENeRF and EvGGS, we use 20 images to fine-tune them and use the rest 180 images for testing. The results are shown in Table.7.
@@ Appendix G Limitations
Although the proposed collaborative learning framework largely improves the performance of the three subtasks, some aspects can still be addressed in the future. First, to accommodate event data, we replaced the spherical harmonics in the original 3DGS with a single intensity value, which reduces its capability to model view-dependent effects. This might be solved by displaying and modeling the lighting direction through a separate network pathway. Second, due to the limitations of 3DGS, this method cannot effectively reconstruct specular metals, such as shown in Fig. (b). We believe one can integrate the idea in Ref-NeRF (Verbin et al., 2022) into 3DGS to address this challenge in the future.

### v3

@@ Abstract
Event cameras offer promising advantages such as high dynamic range and low latency, making them well-suited for challenging lighting conditions and fast-moving scenarios. However, reconstructing 3D scenes from raw event streams is difficult because event data is sparse and does not carry absolute color information. To release its potential in 3D reconstruction, we propose the first event-based generalizable 3D reconstruction framework, called EvGGS, which reconstructs scenes as 3D Gaussians from only event input in a feedforward manner and can generalize to unseen cases without any retraining. This framework includes a depth estimation module, an intensity reconstruction module, and a Gaussian regression module. These submodules connect in a cascading manner, and we collaboratively train them with a designed joint loss to make them mutually promote. To facilitate related studies, we build a novel event-based 3D dataset with various material objects and calibrated labels of grayscale images, depth maps, camera poses, and silhouettes. Experiments show models that have jointly trained significantly outperform those trained individually. Our approach performs better than all baselines in reconstruction quality, and depth/intensity predictions with satisfactory rendering speed. Code and Dataset are demonstrated https://github.com/Mercerai/EvGGS/
@@ 1 Introduction
The contributions of this paper are summarized as follows: (1) We first propose the pure event-based, generalizable 3DGS framework (EvGGS), which faithfully reconstructs 3D scenes as 3D Gaussians from raw event streams and generalizes to various unseen scenarios. The proposed method outperforms existing event-based methods.
@@ 5.1 Event-based 3D Dataset
Dataset Existing event-based 3D datasets such as (Rudnev et al., 2023; Zhou et al., 2018) only contain a limited number of objects and lack high-quality intensity, depth, and mask groundtruths because they mainly concentrate on single scene reconstruction or sparse vision tasks. To fill the current gaps in the community, we establish a full event-based 3D dataset including completed labels, referred to as Ev3DS. The dataset includes a wide variety range of materials. There are 64 objects for training and 15 objects The dataset is constructed and rendered via Blender, it encompasses a multitude of photo-realistic objects, characterized by their complex and varied geometric structures and texture information. We have harnessed VisionBlender(Cartucho et al., 2020) to gather dense depth, mask, and pose information. In each scene, events are generated by a virtual event camera that orbits around the origin of the object in space. We employ V2E(Hu et al., 2021) to generate synthetic event streams maintaining default noise configurations. Additionally, to verify the robustness of the proposed method, we establish and release a novel realistic dataset utilizing the event camera DVXplore for further faithful evaluations, referred to as Ev3D-R. The real-world dataset is essential because the real event camera will raise noises and be more sensitive to illumination changes than synthetic event data. We evaluate the scalability and generalization of the proposed methods on the realistic event data. A detailed introduction to Ev3D-R can be found in the Appendix.
@@ 5.2 Performance of Neural Reconstruction
Table 1 reports all metrics to quantitatively compare the results, which indicates that our method achieves the best performance. Gen. means whether the corresponding methods can generalize to unobserved scenes. Real-time means the corresponding methods can render several images within 1 second. More analysis of this point is in the Appendix. Fig. 2 presents the qualitative results.
@@ 5.4 Quality of Depth Estimation
5.5 Performance on Realistic Event Data
@@ 5.6 Ablation Studies
This subsection demonstrates the impact of different training strategies on model performance in the tasks of novel view synthesis and depth estimation. w/o Joint denotes that we only train the Gaussian regressor with the individually trained and frozen depth estimator and intensity reconstructor. w/o Cascade means that the input does not contain the feature map of the previous network, but only the event voxel and prediction results from the last modules. In Table. 4, the performance of w/o Joint and w/o Cascade is significantly degraded because submodules hardly benefit from the others in the two settings. Besides, w/o ${L}_{I}$ and w/o ${L}_{D}$ represent the corresponding loss variants by removing $L_{I_{\theta}}$ and $L_{D_{\phi}}$ respectively. Table. 4 demonstrates that the absence of depth supervision during joint training leads to a decline in depth estimation. The w/o $L_{I}$ results in a lack of constraints for intensity reconstruction, which degrades the performance of the subsequent cascaded Gaussian regressor and adversely affects the other two submodules with varying degrees.
@@ 6 Conclusion
Acknowledgements
This work is supported by the Guangzhou-HKUST(GZ) Joint Funding Program under Grant No. 2023A03J0682.
Impact Statement
This work aims to advance the pure event-based 3D reconstruction to release the potential of the brain-inspired camera including high dynamic range, low temporal latency, etc. This work is the first to use only event data to reconstruct 3D Gaussians in a generalizable way.
@@ Appendix F Additional Experiments
F.1 The Data Collection Pipeline to Obtain The Realistic Dataset: EvGGS-R
In this subsection, We briefly introduce our data collection pipeline. This realistic dataset is captured by the DVXplore event camera. First, we render RGB images and generate corresponding depth groundtruth via Blender, this step is similar to the synthetic dataset. Then we display the videos for these objects on a high fresh rate screen in our work studio with constant low lighting conditions. Meanwhile, we use the DVXplore to continuously capture the screen to obtain the corresponding realistic event stream. Before collecting data, we have already well-calibrated these devices to ensure high-level data association by using chessboard calibration and image, depth, and event frame alignment. By using the data collection pipeline, we can obtain a realistic event dataset with corresponding images, depths, and 3D model labels as well. Moreover, The distribution of events captured by the real DVXplore will be more complex, realistic, and disordered, which places higher demands on models.
@@ F.2 Qualitative Experiments on Large Scale Scenes
We compare our method with GrayNeRF and E-ENeRF. GrayNeRF means we trained the original NeRF by the grayscale images converted from the RGB images. We optimize the GrayNeRF from scratch because it is a per-scene optimization method.
@@ Appendix G Limitations
Although the proposed collaborative learning framework largely improves the performance of the three subtasks, some aspects can still be addressed in the future. First, to accommodate event data, we replaced the spherical harmonics in the original 3DGS with a single intensity value, which reduces its capability to model view-dependent effects. This might be solved by displaying and modeling the lighting direction through a separate network pathway. Second, this method cannot effectively reconstruct specular metals.